OpenStudy (anonymous):

x/6 - 7 = 2 there is something im not doing right pls help me out i need the layout not the answer

OpenStudy (mrnood):

to solve this you need to get all the terms with x on one side of the equals sign, and all the constants on the other so starting with \[\frac{ x }{ 6 } -7 =2\] start by ADDING 7 to each side write the new equation here...

OpenStudy (anonymous):

x/6 = 9

OpenStudy (mrnood):

ok good so now you have x/6 but you want x so multiply both sides by 6...

OpenStudy (anonymous):

9/6 is where im confused

OpenStudy (anonymous):

oh multiply ok i get it

OpenStudy (anonymous):

x = 54

OpenStudy (mrnood):

that's correct - well done

OpenStudy (anonymous):

thx 4 giving me the layout man

OpenStudy (mrnood):

you should check it yourself by putting 54 into the original equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

54/6 - 7 9 - 7 6

OpenStudy (mrnood):

54/6 - 7 = 9-7 =2 so correct...